OpenStudy (anonymous):

4x2 + 7x + 1 = 0 solve using the quadratic formula

OpenStudy (anonymous):

what was the quadratic formula again?

OpenStudy (anonymous):

a quadratic equation is any equation having the form where x represents an unknown, and a, b, and c are constants with a not equal to 0. If a = 0, then the equation is linear, not quadratic.

OpenStudy (anonymous):

so a= 4 b= 7 c = 1

OpenStudy (anonymous):

\[\frac{ -b \frac{ + }{ - }\sqrt{b^2 - 4ac} }{ 2a} \]

OpenStudy (anonymous):

now that you know the quadratic formula and a,b and c, you can simply plug it in

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

\[\frac{ -7\frac{ + }{ - } \sqrt{16-4(4)(1)}}{ 2(4) }\]

OpenStudy (anonymous):

thank you! i got it right!

OpenStudy (anonymous):

Happy to hear that!